OpenStudy (anonymous):

the integral of (x/(sq root (16-x^2)) dx

OpenStudy (anonymous):

try \(u=16-x^2, du =-2xdx\) should work nicely

OpenStudy (anonymous):

first use substitution then u=x^(2) du/2 = xdx Then use trig substution u = sin(theta)